solve the system by addition or subtitution.
3x-4y=8
6x-2y=10

Answer by checkley75(3666)

MULTIPLY EQUATION 1 BY -2 & THEN ADD:
-6X+8Y=-16
6X-2Y=10
-------------------
6Y=-6
Y=-6/6
Y=-1 ANSWER.
3X-4*-1=8
3X+4=8
3X=8-4
3Y=4
X=4/3 ANSWER.
PROOF
6*4/3-2*-1=10
24/3+2=10
8+2=10
10=10